# Building Age Estimation: A New Multi-Modal Benchmark Dataset and Community Challenge

### v1

@@ V Evaluation and Results
Evaluation metrics. In this work, for the numerical evaluation results, we compute the classification confusion matrix for the task of predicting the construction period of buildings, where we have $7$ different classes, for example “After 2006”. We calculate the average of the diagonal items of the confusion matrix, and we report this metric, which is also known as the Mean Producer Accuracy (MPA) evaluation metric.
@@
We present the evaluation of the models from the challenge in Figs. 11 and 12. We use MPA_ALL_MODALITIES to denote the setting when both training and testing are performed with all modalities. In addition, we also use MPA_TOP_VIEW to denote the setting when training is performed with all modalities and inference is performed using only the two top-view modalities. TOTAL is the average of the two MPAs.
@@
In Figs. 11 and 12, in the results [^2] https://platform.ai4eo.eu/map-your-city/leaderboard, the MPA_TOP_VIEW scores are lower than MPA_ALL_MODALITIES. The top score for MPA_ALL_MODALITIES in the private leaderboard is $0.74$ indicating that the estimation of the construction epoch of buildings in unseen cities is possible. When using only the top-view data during inference, then the performance of the models decreases by more than $10$ points in the MPA score. Finally, we observe that for the TOTAL scores, the top winning values are approximately $0.65$ in the private leaderboard.
VI Conclusion
@@ Acknowledgment
The challenge was accomplished thanks to the contributions of many people. We want to thank all the team: Juan Pedro, Dennis Albrecht, and everyone in ESA AI4EO.

### v2

@@ X Evaluation and Results
In this work, for the numerical evaluation results, we compute the classification confusion matrix for the task of predicting the construction period of buildings, where we have seven different classes, for example “After 2006”. We calculate the average of the diagonal items of the confusion matrix, and we report this metric, which is also known as the Mean Producer Accuracy (MPA) evaluation metric.
@@
We present the evaluation results of the models from the challenge in Figs. 6 and 7. We use MPA_ALL_MODALITIES to denote the setting when both training and testing are performed with all modalities. In addition, we also use MPA_TOP_VIEW to denote the setting when training is performed with all modalities and inference is performed using only the two top-view modalities. TOTAL is the average of the two MPAs.
@@
In Figs. 6 and 7, in the results [^2] https://platform.ai4eo.eu/map-your-city/leaderboard, the MPA_TOP_VIEW scores are lower than MPA_ALL_MODALITIES. The top score for MPA_ALL_MODALITIES in the private leaderboard is $0.74$ indicating that the estimation of the construction epoch of buildings in unseen cities is possible using these models. Furthermore, when using only the top-view data during inference, then the performance of the models decreases by more than $10$ points in the MPA score. Finally, we observe that for the TOTAL scores, the top winning values are approximately $0.65$ in the private leaderboard.
XI Conclusion
@@ Acknowledgment
The challenge was accomplished thanks to the contributions of many people. We want to thank all the team: Juan Pedro, Dennis Albrecht, and everyone in ESA AI4EO http://ai4eo.eu.